OpenStudy (anonymous):

find the equation of the line with the slope 8 passing through (-5,-6)

OpenStudy (anonymous):

do u know the general equation of the slope?

OpenStudy (abb0t):

\((y-y_1)=m(x-x_1) \) or \(y = mx+b\) where m= slope and y = y-coordinate

OpenStudy (anonymous):

exactly.. what @abb0t said. what do u think you'll need to do now?

OpenStudy (anonymous):

exactly.. what @abb0t said. what do u think you'll need to do now?